OpenStudy (anonymous):

y varies inversely with twice the value of x. When y = 8, x = 2. Find y when x = 8.

OpenStudy (anonymous):

Inverse Formula: y=k/x So, 8=k/2 You Mulitply 8 by 2, and you get 16. K=16 . y=16/8 (You Divide by 8, Since x=8) And Your answer will be 2!!!!